The US national debt at the beginning of 2008 was about $9.5 trillion, and at the beginning of 2010 it was about 12.8 trillion. Write a linear function, d, to predict the national debt for any year, x. Then predict the debt at the beginning of 2020. Explain why you think your linear function is or is not a good model for the national debt.
Answer by Boreal(15235)

Let 2008=0
2010=2
so (0, 9.5) and (2, 12.8)
slope is 1.65
point slope formula y-y1=m(x-x1), m slope (x1, y1) point
y-9.5=1.65x
y=1.65x+$9.5 in trillions
2020 is where x=12
y=3.3+1.65x
=9.5+$19.8
=$29.3 trillion.
-
Can simply plot the graph against the actual debt and see what it looks like. There are many non-linear things such as interest accumulation, and there are other variables that may average out to be linear, but it is not clear. There are enough data available to plot and look at the validity of the model.